# Jammer Mitigation in Absorptive RIS-Assisted Uplink NOMA

### v1

@@ V-B Performance for Various A-RIS Locations
TABLE II: Summary of Deployment Parameters with Moving A-RIS.
@@
We adopt a two-dimensional deployment model described by the ($x$, $y$) positions of the nodes, as defined in Fig. 9. We then position the A-RIS at various locations along the $x$-axis, with the BS at the origin, and the jammer and users at the fixed locations given in Table II. The related path losses are then calculated for all the channel components.
@@
VI Discussion and Future Work
VI-A Relevance of the Results for Downlink Scenario
In this section, we discuss the relevance of the results above for a corresponding downlink scenario, as illustrated in Fig. 12. We discuss an initial system model as a starting point for possible future work on A-RIS optimization. In general, the downlink problem is quite different, since in the downlink the jammer has more than one potential target, more precisely all the $K$ users. Thus, to analyze the downlink problem, we would need to define the objective of the jammer:
@@ VI-A Relevance of the Results for Downlink Scenario
In Case 1, the jammer will optimize its signal focusing on attacking only a selected victim receiver. This case would be similar to the uplink case, in which the jammer focuses on optimizing its attack only on the BS. However, a key difference is that in our uplink model, there is only one receiver, the BS, whereas in the downlink there are still $K-1$ other active receivers besides the one under attack that would still experience the jammer interference. Without loss of generality, assume that the jammer attacks user $k_{v}$. Furthermore, assume a typical power domain downlink NOMA system, in which the users’ path losses are ordered with the strongest channel from user $k=1$, and the weakest channel for the last user $k=K$. Then, unless the target downlink SINRs $T_{k}$ are very different for the users, the BS will allocate more power to users with weaker channels. This enables each user to perform SIC for all other users with weaker path loss, i.e. higher user index $k$, before decoding its own signal. Under these assumptions, and assuming equal AWGN power $\sigma^{2}$ for all users, with perfect SIC the resulting SINR expressions become as follows, where $\sigma_{\text{j}}^{2}(k_{v},k)$ denotes the jammer power at user $k$ optimized to attack user $k_{v}$:
$\gamma_{1}=\frac{p_{1}|h_{1}+{\bf g}_{1}^{\top}\boldsymbol{\Phi}{\bf f}|^{2}}{\sigma_{\text{j}}^{2}(k_{v},1)+\sigma^{2}},$ (21)
$\gamma_{2}=\frac{p_{2}|h_{2}+{\bf g}_{2}^{\top}\boldsymbol{\Phi}{\bf f}|^{2}}{p_{1}|h_{2}+{\bf g}_{2}^{\top}\boldsymbol{\Phi}{\bf f}|^{2}+\sigma_{\text{j}}^{2}(k_{v},2)+\sigma^{2}},$ (22)
$\gamma_{3}=\frac{p_{3}|h_{3}+{\bf g}_{3}^{\top}\boldsymbol{\Phi}{\bf f}|^{2}}{(p_{1}+p_{2})|h_{3}+{\bf g}_{3}^{\top}\boldsymbol{\Phi}{\bf f}|^{2}+\sigma_{\text{j}}^{2}(k_{v},3)+\sigma^{2}},$ (23)
and in general for user $k=1,\,\ldots,\,K$
$\gamma_{k}=\frac{p_{k}|h_{k}+{\bf g}_{k}^{\top}\boldsymbol{\Phi}{\bf f}|^{2}}{\smashoperator[]{\sum_{}^{}}_{i=1}^{k-1}p_{i}|h_{k}+{\bf g}_{k}^{\top}\boldsymbol{\Phi}{\bf f}|^{2}+\sigma_{\text{j}}^{2}(k_{v},k)+\sigma^{2}}.$ (24)
In Case 1, with a single-antenna victim receiver, the intelligent jammer would optimize its jamming signal ${\bf x}_{\text{j}}$ as in the uplink case described in Section II-B. Thus, it is straightforward to obtain the expression for the resulting received jammer power at user $k$ optimized for jamming user $k_{v}$, as follows.
$\sigma_{\text{j}}^{2}(k_{v},k)=P_{\text{j}}\frac{\|({\bf h}_{\text{j},k}^{\top}+{\bf g}_{k}^{\top}\boldsymbol{\Phi}{\bf G}_{\text{j}})({\bf h}_{\text{j},k_{v}}^{\top}+{\bf g}_{k_{v}}^{\top}\boldsymbol{\Phi}{\bf G}_{\text{j}})^{H}\|^{2}}{\|{\bf h}_{\text{j},k_{v}}^{\top}+{\bf g}_{k_{v}}^{\top}\boldsymbol{\Phi}{\bf G}_{\text{j}}\|^{2}}\;,$ (25)
which for the victim UE $k_{v}$ becomes
$\sigma_{\text{j}}^{2}(k_{v},k_{v})=P_{\text{j}}\|{\bf h}_{\text{j},k_{v}}^{\top}+{\bf g}_{k_{v}}^{\top}{\boldsymbol{\Phi}}{\bf G}_{\text{j}}\|^{2}\;,~{}k_{v}=1,\,\ldots,\,K.$ (26)
This is an expression similar to the one when the jammer attacks the BS in the uplink case. If we would like to optimize the total sum transmit power in this downlink case, we observe that similar to the uplink case, $\gamma_{1}$ should equal $T_{1}$, allowing us to eliminate one optimization parameter as in (9) (in the uplink we eliminate $p_{K}$). However, since in general the path losses $\|{\bf h}_{\text{j},k}\|$ are different between the jammer and the $K$ different users, there is no straightforward way to map the power minimization problem in the downlink to Algorithm 1 due to point-to-multipoint channels from the jammer to the users, resulting in a user index $k$-dependent jammer power $\sigma_{\text{j}}^{2}(k_{v},k)$.
@@
In Case 2, one could define the jamming signal as a weighted sum of the jamming signals targeting each user, i.e. $\bar{\bf x}_{\text{j}}=(\sqrt{P}_{\text{j}}/{\bar{\rho}})\left(\bar{\bf h}_{\text{j}}^{\top}+\bar{\bf g}^{\top}\boldsymbol{\Phi}{\bf G}_{\text{j}}\right)^{H}$, where $\bar{\rho}=\|\bar{\bf h}_{\text{j}}^{\top}+{\bar{\bf g}}^{\top}{\boldsymbol{\Phi}}{\bf G}_{\text{j}}\|$, $\bar{\bf h}_{\text{j}}=\sum_{i=1}^{K}\eta_{i}{\bf h}_{\text{j},i}$, and $\bar{\bf g}=\sum_{i=1}^{K}\eta_{i}{\bf g}_{i}$, for some user weights $\{\eta_{k}\},k=1,\ldots K$ chosen by the jammer. In that case, it is straightforward to show that the jamming power at user $k$ depends only on $k$,
$\sigma_{\text{j}}^{2}(k)=P_{\text{j}}\frac{\|({\bf h}_{\text{j},k}^{\top}+{\bf g}_{k}^{\top}\boldsymbol{\Phi}{\bf G}_{\text{j}})(\bar{\bf h}_{\text{j}}^{\top}+\bar{\bf g}^{\top}\boldsymbol{\Phi}{\bf G}_{\text{j}})^{H}\|^{2}}{\|\bar{\bf h}_{\text{j}}^{\top}+\bar{\bf g}^{\top}\boldsymbol{\Phi}{\bf G}_{\text{j}}\|^{2}}\;.$ (27)
@@
In Case 3, one could define the objective of the jammer to maximize $\sigma_{\text{j}}^{2}(k_{v},k)=\sigma_{j,0}^{2}$, i.e., a constant received jammer power for all users, independent of $k$. For propagation conditions in which there would be a feasible solution for the jammer to meet this objective, we would end up with an optimization problem that could potentially be solved using Algorithm 1.
VI-B Additional A-RIS Considerations
In Section III we discussed a potential implementation of an A-RIS and the resulting non-ideal amplitude/phase coverage of the resulting A-RIS reflection coefficients. In particular, we defined $\beta_{\mathrm{max}}$ as the circle in the complex plane with a maximum radius for which an arbitrary phase response could be achieved. However, there are additional non-ideal A-RIS properties that should be taken into account in practical A-RIS implementations, as outlined below.
  1. 1.
    The UC performance is frequency-dependent. As the frequency spans from microwave to mm-wave, the impact of the aforementioned losses becomes more pronounced. In addition, the quality factor and tuning range of commercial diodes degrade at high frequencies. For instance, due to increased parasitic effects, the total capacitance tuning ratio of the off-the-shelf varactor diode is 3.4 at 10 GHz but drops to only 1.8 at 28 GHz, thus potentially restricting the phase range or $\beta_{\mathrm{max}}$ of the UC [37]. As indicated by reported RIS measurement results, $\beta_{\mathrm{max}}$ achieves 0.7 at 4.15 GHz [38] and 0.5 at 28 GHz under normal plane-wave incidence. Thus, wide-band optimization of realistic A-RIS would need to consider a frequency dependent $\beta_{\mathrm{max}}$.
  2. 2.
    The UC performance is angle-dependent. Note that the UC gain $\beta$ incorporates angular effects, as both $Z_{\mathrm{sub}}$ and $Z_{\mathrm{sh}}$ can vary with both the azimuth and elevation angle of the incident wave. For dipole-like UCs with sub-wavelength sizes, the angular sensitivity is mostly defined by the substrate electrical thickness, whereas $Z_{\mathrm{sh}}$ weakly depends on the incident angle. Typically, the performance of the UC deteriorates at large incident angles, thereby diminishing the coverage area. Thus, for more realistic results above, we should also include a model of the coverage region depending on the azimuth and elevation angle of the incident wave.
  3. 3.
    The UC performance is polarization-dependent. A single-polarized UC is only capable of controlling the reflection coefficient of the co-polarized reflected wave. Full control of the reflection response can be achieved by employing dual-polarized UCs, albeit at the cost of increased design complexity. Independently controlling the reflection response in the polarization domain would have importance in particular for long hops, since the vertical and horizontal polarization components in general experience different path losses. In such cases, an A-RIS model capable of independently controlling the polarization modes and including polarization losses would be important.
VII Conclusion
@@
An optimization problem was formulated whose goal is to minimize the total transmit power of the users under constraints on the SINR that the users achieve at the base station. The optimization is non-convex and of high dimension, so we reformulate it and employ an iterative approach that alternates between using linear programming to estimate the user powers, and the Dinkelbach algorithm to determine the A-RIS coefficients.
Our simulation results demonstrate that A-RIS elements provide dramatic reductions in required total transmit power due to their ability to enhance the signals of interest and cancel the effects of the interference. We also demonstrated that if the number of elements in the intelligent surface is large, the A-RIS can provide additional degrees of freedom that enable increased signal enhancement and interference mitigation. We showed that the influence of the absorption is most important when the ratio of A-RIS elements to jammer antennas is not too large, i.e., when the extra degrees of freedom offered by the variable element amplitudes become crucial.
Finally, we discussed potential future directions for this research: studying the corresponding downlink A-RIS-assisted NOMA problem and considering additional non-ideal A-RIS properties in practical implementations. Further work could also consider the deployment of A-RIS-assisted NOMA for ISAC systems, and study the impact of imperfect SIC on power-domain NOMA assisted by an A-RIS.

### v2

@@ V-B Performance for Various A-RIS Locations
TABLE III: Summary of Deployment Parameters with Moving A-RIS.
@@
We adopt a two-dimensional deployment model described by the ($x$, $y$) positions of the nodes, as defined in Fig. 9. We then position the A-RIS at various locations along the $x$-axis, with the BS at the origin, and the jammer and users at the fixed locations given in Table III. The related path losses are then calculated for all the channel components.
@@
V-C Relevance of the Results for Downlink Scenario
Here we briefly explore a corresponding downlink scenario, as illustrated in Fig. 12, in order to investigate if there is an uplink-downlink duality. In general, the downlink problem is quite different, since the jammer has more than one potential target, more precisely, all the $K$ users. Thus, to analyze the downlink problem, we would need to define the objective of the jammer:
@@ V-C Relevance of the Results for Downlink Scenario
In Case 1, the jammer will optimize its signal, focusing on attacking only a selected victim receiver. This case would be similar to the uplink case, in which the jammer focuses on optimizing its attack only on the BS. However, a key difference is that in our uplink model, there is only one receiver, the BS, whereas in the downlink, there are still K1K-1 other active receivers besides the one under attack that would still experience the jammer interference. In addition, since in general the path losses $\|{\bf h}_{\text{j},k}\|$ are different between the jammer and the $K$ users, there is no straightforward way to map the power minimization problem in the downlink to Algorithm 1 due to point-to-multipoint channels from the jammer to the users, resulting in a user index $k$-dependent jammer power $\sigma_{\text{j}}^{2}(k_{v},k)$, where $k_{v}$ denotes the victim user under jamming attack.
@@
In Case 2, one could define the jamming signal as a weighted sum of the jamming signals targeting each user. In that case, it is straightforward to show that the jamming power at user $k$ would be the same for all victims $k_{v}$, but would still depend on $k$, i.e. $\sigma_{\text{j}}^{2}(k)$.
@@
In Case 3, one could define the objective of the jammer to maximize $\sigma_{\text{j}}^{2}(k_{v},k)=\sigma_{j}^{2}$, i.e., equal received jammer power for all users, independent of $k$. For propagation conditions in which there would be a feasible solution for the jammer to meet this objective, we would end up with an optimization problem that could potentially be solved using Algorithm 1.
VI Conclusion
@@
An optimization problem was formulated whose goal is to minimize the total transmit power of the users under constraints on the SINR that the users achieve at the base station. The optimization is non-convex and of high dimension, so we reformulate it and employ an iterative approach that alternates between using linear programming to estimate the user powers and the Dinkelbach algorithm to determine the A-RIS coefficients.
Our simulation results demonstrate that A-RIS elements provide dramatic reductions in required total transmit power due to their ability to enhance the signals of interest and cancel the effects of the interference. We also demonstrated that if the number of elements in the intelligent surface is large, the A-RIS can provide additional degrees of freedom that enable increased signal enhancement and interference mitigation. We showed that the influence of the absorption is most important when the ratio of A-RIS elements to jammer antennas is not too large, i.e., when the extra degrees of freedom offered by the variable element amplitudes become crucial. Finally, we briefly discussed the relevance of our results for a corresponding downlink scenario.